# Boolean Nearest Neighbor Language in the Knowledge Compilation Map

### v1

@@ 6 Conclusions
Next, we have studied the complexity status of standard transformations and queries for the BNN language. Although it supports a decent subset of queries in polynomial time (CO, VA, IM, CE, ME) and hence it passes the necessary condition for a target compilation language formulated in [10] [^1] Here we refer to the following sentence from [10]: ”For a language to qualify as a target compilation language we will require that it permits a polytime clausal entailment test.”, we feel that the lack of supported transformations (only negation is supported), and in particular the fact that conditioning is not supported in polynomial time, disqualifies the BNN from being a good target language for knowledge compilation. The open problems left for future research is the complexity staus of the EQ, SE, and CT queries and of the SFO transformation.
@@ 7 Acknowledgement
We are grateful to Petr Kučera for several helpful suggestions and for the example that is used in the proof of Theorem 4.5.

### v2

@@ 6 Conclusions
Next, we have studied the complexity status of standard transformations and queries for the BNN language. Although it supports a decent subset of queries in polynomial time (CO, VA, IM, CE, ME) and hence it passes the necessary condition for a target compilation language formulated in [10] [^2] Here we refer to the following sentence from [10]: ”For a language to qualify as a target compilation language we will require that it permits a polytime clausal entailment test.”, we feel that the lack of supported transformations (only negation is supported), and in particular the fact that conditioning is not supported in polynomial time, disqualifies the BNN from being a good target language for knowledge compilation. The open problem left for future research is the complexity status of the singleton forgetting transformation.
@@ 7 Acknowledgement
We are grateful to Petr Kučera for several helpful suggestions and for the construction that is used in the proof of Theorem 4.5.